In triangle ABC, linr CD is both the median and the altitude to side AB. If AB=5x+3, AC=2x=8, and BC=3x+5, what is the perimeter of triangle ABC?
Answer by ankor@dixie-net.com(12692)

In triangle ABC, line CD is both the median and the altitude to side AB. If AB=5x+3, AC=2x+8, and BC=3x+5, what is the perimeter of triangle ABC?
:
If the median is the same as the altitude, this triangle has two equal sides,
namely, AC and BC, therefore:
2x + 8 = 3x + 5
8 - 5 = 3x - 2x
x = 3
:
AB: 5(3) + 3 = 18
AC: 2(3) + 8 = 14
BC: 3(3) + 5 = 14
------------------
the perimeter: 46